# M+: Extending MemoryLLM with Scalable Long-Term Memory

### v1

@@ 2.1 Token-level Memory
(a) Update
(b) Generation
Figure 1: Update and Generation Process of MemoryLLM. We process the chunk with $\phi_{l}$ to obtain new $K$ tokens during the update process, which is perceived by $\phi$ using cross-attention during the generation process.
(a) Update Process.
(b) Generation
Figure 2: Update and Generation Process of M+. For layer $l$, during Update, the old memory pool $\theta_{l}$ is split into two parts: $K$ dropped tokens and $N-K$ remaining tokens. The dropped tokens are stored in the long-term memory $\Theta_{l}$ while the remaining tokens and new $K$ tokens are combined to obtain the new memory pool $\theta_{l}^{\prime}$. Then during generation, we use our co-trained retriever to retrieve tokens from $\Theta_{l}$, which is fed into the transformer layer $\phi_{l}$ along with the short-term memory $\theta_{l}$ and the query hidden states.
@@ 2.2 Latent-Space Memory
Latent-space memory stores information in a compressed format, embedding knowledge into hidden states (Khandelwal et al., 2019; Wang et al., 2024a), model parameters (Wang et al., 2024c), or an external latent space (Das et al., 2024), among other methods. Some approaches use memory slots to encode information (Al Adel & Burtsev, 2021), while others rely on key-value caches stored in memory pools for future retrieval (Wu et al., 2022b; Wang et al., 2023; He et al., 2024; Park & Bak, 2024). Notably, CamelLoT (He et al., 2024) and Memoria (Park & Bak, 2024) incorporate forgetting mechanisms to better emulate human memory. Similarly, MemoryLLM (Wang et al., 2024a) compresses knowledge into hidden states and employs random dropping to prevent unbounded memory growth. The M3 method (Yang et al., 2024) also stores memory in the hidden-state space, archiving a vast pretraining dataset comprising $1.1\times 10^{8}$ text chunks. Distinct from methods that utilize hidden states or key-value caches, Larimar (Das et al., 2024) introduces a memory matrix that supports read and write operations, demonstrating effectiveness in knowledge-editing tasks. Furthermore, SELF-PARAM (Wang et al., 2024c) explores embedding knowledge directly into model parameters without degrading the model’s capabilities or requiring additional parameters. These latent-space memory techniques have shown promising results across various downstream tasks. By saving information in a compressed format and leveraging retrieval during generation, they enable substantial expansions of the context window without incurring excessive GPU memory costs. Despite the advantages and potential of Latent-Space Memory, existing methods within this category typically fall short when dealing with extremely long input (Das et al., 2024; Wang et al., 2024c, a; He et al., 2024). In contrast, M+ can have much longer retention compared to existing methods.
@@ 3.1 Preliminaries
We first introduce the structure of MemoryLLM (Wang et al., 2024a), which serves as the base structure of M+. MemoryLLM comprises two main components: $\theta$ (the memory pool) and $\phi$ (a transformer-based decoder-only language model). The memory pool $\theta$ consists of $L$ layers: $\{\theta_{l}\}_{l=1}^{L}$ where $L$ is the number of layers in the transformer $\phi$. For every layer, $\theta_{l}$ has $N$ memory tokens, where each token is a vector in $\mathbb{R}^{d}$, with $d$ representing the hidden size of the language model. During the update process, the last $K$ tokens from the $l$-th layer’s memory pool, $\theta_{l}$, are extracted and combined with the chunk to be injected. The resulting new $K$ tokens are then merged back into $\theta_{l}$ (illustrated in Figure 1(a)). Merging is achieved by randomly dropping $K$ tokens from $\theta_{l}$ and appending the new $K$ tokens to the end. During generation, the memory pool $\theta_{l}$ is perceived using cross-attention. (as shown in Figure 1(b)).
@@ 3.2.1 Memory Structures
During the update process, note that in the original MemoryLLM, $K$ tokens are dropped from $\theta$ during updates and are permanently discarded. In M+, the dropped $K$ tokens are instead stored in the long-term memory $\Theta$, ensuring their retention for extended durations (as illustrated in Figure 2(a)). We assign each token the variable “age” so that after retrieving tokens from $\Theta$ we can sort these tokens according to age, ensuring that the tokens are chronologically ordered. As for the new $K$ tokens, they are obtained with the same process as in MemoryLLM, described in Figure 1(a). When the memory tokens in the long-term memory reach the maximum capacity, i.e., $M$ tokens, we would drop the tokens with the largest ages.
@@
In our training, we use two sets of LoRA weights, one is activated during the update process (as shown in Figure 2(a)) and the other is activated during the generation process (as shown in Figure 2(b)). Intuitively, the update process compresses the information (similar to writing) while the generating process loads the information (similar to reading), thus having two LoRA weights could potentially make learning easier for our model. This is similar to the intuition in T5 where they find sharing the weights of encoder and decoder leads to slightly inferior performances (See Table 2 in Raffel et al. (2020)).
@@ 3.2.2 Retriever Design and Training
The retriever has two projectors: query projector $f_{q}$ and key projector $f_{k}$, which are all instantiated with a two-layer perceptron. The output dimension of both projectors, denoted as $d_{proj}$, is set to be a small number. In our experiments, we set $d_{proj}$ to be $d/20$ where $d$ is the hidden size of the language model $\phi$. When dropping tokens from $\theta_{l}$ into $\Theta_{l}$ (as shown in Figure 2(a)), we apply $f_{k}$ on top of the dropped memory tokens, thus we need an additional pool storing all the key vectors corresponding to the memory tokens in $\Theta_{l}$. Note that the key vectors are the output from $f_{k}$, and are of dimension $d_{proj}$, requiring little additional memory footprint compared to the long-term memory. During generation, given the hidden states from the query, we apply $f_{q}$ on the hidden states to get query vectors and use them to retrieve tokens from $\Theta_{l}$ according to the dot product between query vectors and key vectors.
@@ 3.2.3 Training Details
We build M+ on top of Llama-3.1-8B (Dubey et al., 2024) and train it using eight A100 GPUs. We tried FSDP, deepspeed-stage-2, and deepspeed-stage-3 and we finally choose deepspeed-stage-2 due to resource limitation and library incompatibility (see the details in Appendix A). Specifically, we set $K=256$, $N=10240$ ($N$ is the number of tokens in the short-term memory, see Section 3.1), and the number of tokens of extracted LTM in Figure 2(b) is set to 2,560. The generation window (i.e., the maximum length of generation) is set to be 2,048. Thus maximally we have the attention matrix of shape $(12,800+2,048)$ by $2,048$, which is fit into eight A100 GPUs using deepspeed-stage-2. Although Llama-3.1-8B can practically handle a 128k context window, it went through much more extensive training that we cannot afford. Should we have more GPUs and more budget, M+ could also be scaled to 128k level. Given the constraint of GPU resources, we have scaled to 12,800 memory tokens and 2,048 generation context window, relying on Llama-3.1-8B’s capability on a context window of $12,800+2,048=14,848$ tokens. Thus, for fair comparisons, we mainly focus on Llama-3.1-8B-16k as the baseline.
@@ 4.1.2 Experimental Results
The results for both benchmarks are shown in Figure 3. From the results, we observe the following: (1) M+ consistently outperforms all baselines, demonstrating its efficiency and effectiveness. In LongBook-QA, M+ achieves superior QA-F1 scores even while processing the least number of tokens (12,800 tokens in memory and 2,048 tokens in the generation context window). Similarly, in LongBook Event QA, M+ excels at identifying ground-truth events, showcasing its ability to reason over long-term dependencies. (2) Compared to Llama-3.1-3B-128k, the results on dataset Longbook-Event-QA suggest that tasks requiring reasoning capabilities benefit more from larger models with tailored structures for extended context windows rather than smaller models with longer context capacities. This highlights the importance of balancing model size and memory mechanisms under fixed GPU memory budgets. (3) Llama-3.1-8B-SnapKV underperforms Llama-3.1-8B-16k on LongBook-QA, indicating that solely relying on attention scores to select key tokens may not consistently yield optimal results. In contrast, M+ leverages a jointly trained retriever to identify and extract memory tokens, resulting in more effective performance on both datasets. (4) Memory Efficiency: While M+ achieves state-of-the-art results, it does so with a highly memory-efficient design. Detailed memory cost comparisons are provided in Section 4.2.
@@ 4.2 GPU Cost Comparison
In this section, we report the maximum GPU memory allocated during the inference across both datasets for each method mentioned in Section 4.1. The results are shown in Table 1. From the results, we can find that M+ has the lowest GPU memory cost except for Llama-3.1-8B-16k. The reason that M+ use less tokens but cost more GPU is that we have 12,800 tokens in each layer, while Llama-3.1-8B-16k has only one layer of 16k tokens. Therefore, we propose to offload the memory tokens on CPU, and reload them into GPU when the calculation reaches a certain layer. With that, we can sacrifice some I/O time cost but substantially decrease the GPU cost without affecting the performance. This leads to “M+ (offload)” which achieves the least GPU memory consumption.
Table 1: GPU Memory Cost Comparison.
@@ 4.4 Experimental Results on (Relatively) Short Documents
We evaluate the performance of M+ and Llama-3.1-8B on relatively short documents using the LongBench benchmark, considering input lengths of 8k and 16k tokens. The results, presented in Table 2, show that M+ could match the performance of Llama-3.1-8B on 4 out of 6 datasets, except for hotpotqa and musique. This performance difference can be attributed to two primary factors: (1) Random Dropping Mechanism: M+ employs a random dropping mechanism that can lead to partial information loss. For instance, processing an 8k input requires splitting it into 12 chunks (each chunk being 512 tokens), while the last 2k tokens are directly included in the generation context window. The first 6k tokens (12 chunks) are compressed into $2755.6256+256*\frac{N-K}{N}+\cdots+256*(\frac{N-K}{N})^{11}=$ 2755.6 tokens (with around 316.4 memory tokens dropped). For 16k tokens, the first 14k tokens are compressed 5530 tokens (with around 1638 tokens dropped). As some tokens are dropped, the performance may get affected slightly. Note that this could lead to a longer context window while sacrificing some of the performance in relatively shorter context tasks. (2) Limited Cross-Chunk Attention: When processing chunks into memory, M+ uses the last $K$ tokens and the hidden states from the new chunk as input (illustrated in Figur e2(a)). In contrast, Llama-3.1-8B processes each chunk with access to all previous tokens, enabling cross-attention between chunks. While this approach allows Llama-3.1-8B to maintain full attention across chunks, it comes with significantly higher computational and memory costs due to the quadratic scaling of transformer computations. In comparison, M+ achieves linear computational scaling, making it more GPU-memory-efficient for processing extremely long inputs (see Section 4.2), albeit with some trade-off in performance in tasks with relatively shorter contexts.
@@ 4.6.1 Retrieval Quality
In our implementation, the long-term memory is initially of size 5120, and then it gradually increases to 80k in our knowledge retention experiments (it hits 81,276 when there 160k tokens are injected). To access retrieval quality, we leverage the knowledge retention task with SQuAD dataset, where the first $K=256$ tokens are critical for answering the questions. These $K=256$ tokens are denoted as ground-truth tokens. We track the number of ground-truth tokens in the long-term memory and how many tokens are retrieved back into the “Extracted LTM” pool in Figure 2(b) when queried after various numbers of tokens are injected. We present the results in Figure 7, demonstrating the retrieval quality as more tokens are dropped from the memory pool to the long-term memory. From the figure we can see that around 30% tokens are retrieved. For reference, random retrieval would retrieve $2,560/81,276=3\%$ tokens.

### v2

@@ 2.1 Token-level Memory
Figure 1: The left side shows the Update and Generation Process of MemoryLLM (Wang et al., 2024a). We process the chunk with $\phi_{l}$ to obtain new $K$ tokens during the update process, which is perceived by $\phi$ using cross-attention during the generation process. The right side shows the Update and Generation Process of M+. For layer $l$, during Update, the old memory pool $\theta_{l}$ is split into two parts: $K$ dropped tokens and $N-K$ remaining tokens. The dropped tokens are stored in the long-term memory $\Theta_{l}$ while the remaining tokens and new $K$ tokens are combined to obtain the new memory pool $\theta_{l}^{\prime}$. Then during generation, we use our co-trained retriever to retrieve tokens from $\Theta_{l}$, which is fed into the transformer layer $\phi_{l}$ along with the short-term memory $\theta_{l}$ and the query hidden states. The major difference between MemoryLLM and M+ is the introduction of Long-Term Memory $\Theta_{l}$.
@@ 2.2 Latent-Space Memory
Latent-space memory stores information in a compressed format, embedding knowledge into soft prompts (Rakotonirina & Baroni, 2024), hidden states (Khandelwal et al., 2019; Bulatov et al., 2022, 2023; Wang et al., 2024a), model parameters (Wang et al., 2024c), or an external latent space (Das et al., 2024), among other methods. Some approaches use memory slots to encode information (Al Adel & Burtsev, 2021), while others rely on key-value caches stored in memory pools for future retrieval (Wu et al., 2022b; Wang et al., 2023; He et al., 2024; Park & Bak, 2024). Notably, CamelLoT (He et al., 2024) and Memoria (Park & Bak, 2024) incorporate forgetting mechanisms to better emulate human memory. Similarly, MemoryLLM (Wang et al., 2024a) compresses knowledge into hidden states and employs random dropping to prevent unbounded memory growth. The M3 method (Yang et al., 2024) also stores memory in the hidden-state space, archiving a vast pretraining dataset comprising $1.1\times 10^{8}$ text chunks. Distinct from methods that utilize hidden states or key-value caches, Larimar (Das et al., 2024) introduces a memory matrix that supports read and write operations, demonstrating effectiveness in knowledge-editing tasks. Furthermore, SELF-PARAM (Wang et al., 2024c) explores embedding knowledge directly into model parameters without degrading the model’s capabilities or requiring additional parameters. These latent-space memory techniques have shown promising results across various downstream tasks. By saving information in a compressed format and leveraging retrieval during generation, they enable substantial expansions of the context window without incurring excessive GPU memory costs. Despite the advantages and potential of Latent-Space Memory, existing methods within this category typically fall short when dealing with extremely long input (Das et al., 2024; Wang et al., 2024c, a; He et al., 2024). In contrast, M+ can have much longer retention compared to existing methods.
@@ 3.1 Preliminaries
We first introduce the structure of MemoryLLM (Wang et al., 2024a), which serves as the base structure of M+. MemoryLLM comprises two main components: $\theta$ (the memory pool) and $\phi$ (a transformer-based decoder-only language model). The memory pool $\theta$ consists of $L$ layers: $\{\theta_{l}\}_{l=1}^{L}$ where $L$ is the number of layers in the transformer $\phi$. For every layer, $\theta_{l}$ has $N$ memory tokens, where each token is a vector in $\mathbb{R}^{d}$, with $d$ representing the hidden size of the language model. During the update process, the last $K$ tokens from the $l$-th layer’s memory pool, $\theta_{l}$, are extracted and combined with the chunk to be injected. The resulting new $K$ tokens are then merged back into $\theta_{l}$ (illustrated in Figure 1). Merging is achieved by randomly dropping $K$ tokens from $\theta_{l}$ and appending the new $K$ tokens to the end. During generation, the memory pool $\theta_{l}$ is perceived using cross-attention.
@@ 3.2.1 Memory Structures
During the update process, note that in the original MemoryLLM, $K$ tokens are dropped from $\theta$ during updates and are permanently discarded. In M+, the dropped $K$ tokens are instead stored in the long-term memory $\Theta$, ensuring their retention for extended durations (as illustrated in Figure 1). We assign each token the variable “age” so that after retrieving tokens from $\Theta$ we can sort these tokens according to age, ensuring that the tokens are chronologically ordered. As for the new $K$ tokens, they are obtained with the same process as in MemoryLLM, described in Figure 1. When the memory tokens in the long-term memory reach the maximum capacity, i.e., $M$ tokens, we would drop the tokens with the largest ages.
@@
In our training, we use two sets of LoRA weights, one is activated during the update process, and the other is activated during the generation process (as shown in Figure 1). Intuitively, the update process compresses the information (similar to writing) while the generating process loads the information (similar to reading), thus having two LoRA weights could potentially make learning easier for our model. This is similar to the intuition in T5 where they find sharing the weights of encoder and decoder leads to slightly inferior performances (See Table 2 in Raffel et al. (2020)).
@@ 3.2.2 Retriever Design and Training
The retriever has two projectors: query projector $f_{q}$ and key projector $f_{k}$, which are all instantiated with a two-layer perceptron. The output dimension of both projectors, denoted as $d_{proj}$, is set to be a small number. In our experiments, we set $d_{proj}$ to be $d/20$ where $d$ is the hidden size of the language model $\phi$. When dropping tokens from $\theta_{l}$ into $\Theta_{l}$ (as shown in Figure 1), we apply $f_{k}$ on top of the dropped memory tokens, thus we need an additional pool storing all the key vectors corresponding to the memory tokens in $\Theta_{l}$. Note that the key vectors are the output from $f_{k}$, and are of dimension $d_{proj}$, requiring little additional memory footprint compared to the long-term memory. During generation, given the hidden states from the query, we apply $f_{q}$ on the hidden states to get query vectors and use them to retrieve tokens from $\Theta_{l}$ according to the dot product between query vectors and key vectors.
@@ 3.2.3 Training Details
We build M+ on top of Llama-3.1-8B (Dubey et al., 2024) and train it using eight A100 GPUs. We tried FSDP, deepspeed-stage-2, and deepspeed-stage-3 and we finally choose deepspeed-stage-2 due to resource limitation and library incompatibility (see the details in Appendix A). Specifically, we set $K=256$, $N=10240$ ($N$ is the number of tokens in the short-term memory, see Section 3.1), and the number of tokens of extracted LTM in Figure 1 is set to 2,560. The generation window (i.e., the maximum length of generation) is set to be 2,048. Thus maximally we have the attention matrix of shape $(12,800+2,048)$ by $2,048$, which is fit into eight A100 GPUs using deepspeed-stage-2. Although Llama-3.1-8B can practically handle a 128k context window, it went through much more extensive training that we cannot afford. Should we have more GPUs and more budget, M+ could also be scaled to 128k level. Given the constraint of GPU resources, we have scaled to 12,800 memory tokens and 2,048 generation context window, relying on Llama-3.1-8B’s capability on a context window of $12,800+2,048=14,848$ tokens. Thus, for fair comparisons, we mainly focus on Llama-3.1-8B-16k as the baseline.
@@ 4.1.1 Experimental Settings
(1) Llama-3.1-3B-128k: A 3B parameter version of the Llama3 series. This model uses a 128k context window and consumes comparable GPU memory to M+ because of its smaller size. (2) Llama-3.1-8B-16k: The original model with context window fixed as 16k. (3) Llama-3.1-8B-BM25: We use BM25 as the retriever. Specifically, we process the entire book by dividing it into chunks of 4,096 tokens and retrieve four relevant chunks for each question. The model Llama-3.1-8B is required to answer the question with four retrieved chunks. (4) Llama-3.1-8B-SnapKV, We processes a 32k token input and dynamically selects 16k key-value caches from the saved 32k caches with the techniques introduced from SnapKV (Li et al., 2024). SnapKV incurs significant memory overhead, as illustrated in Section 4.2.
We present the primary results in the main paper, while deferring extended discussions and supplementary experiments to the appendix. These include: similarities to attention-based retrieval methods (Appendix E.1); the structure of long-term memory (Appendix E.2); latency and memory consumption when scaling (Appendix E.3); FLOPs comparison (Appendix E.4); and the interpretability of memory vectors (Appendix E.5).
@@ 4.1.2 Experimental Results
The results for both benchmarks are shown in Figure 2. From the results, we observe the following: (1) M+ consistently outperforms all baselines, demonstrating its efficiency and effectiveness. In LongBook-QA, M+ achieves superior QA-F1 scores even while processing the least number of tokens (12,800 tokens in memory and 2,048 tokens in the generation context window). Similarly, in LongBook Event QA, M+ excels at identifying ground-truth events, showcasing its ability to reason over long-term dependencies. (2) Compared to Llama-3.1-3B-128k, the results on dataset Longbook-Event-QA suggest that tasks requiring reasoning capabilities benefit more from larger models with tailored structures for extended context windows rather than smaller models with longer context capacities. This highlights the importance of balancing model size and memory mechanisms under fixed GPU memory budgets. (3) Llama-3.1-8B-SnapKV underperforms Llama-3.1-8B-16k on LongBook-QA, indicating that solely relying on attention scores to select key tokens may not consistently yield optimal results. In contrast, M+ leverages a jointly trained retriever to identify and extract memory tokens, resulting in more effective performance on both datasets. (4) M+ outperforms Llama-3.1-8B with BM25 retriever, and Llama-3.1-8B with BM25 does not consistently outperform the original model Llama-3.1-8B-16k, particularly in tasks like Longbook-Event-QA which requires a more global understanding of the entire narrative. This highlights the limitations of chunk-level retrieval in scenarios that demand long-range comprehension. (5) Memory Efficiency: While M+ achieves state-of-the-art results, it does so with a highly memory-efficient design. Detailed memory cost comparisons are provided in Section 4.2.
@@ 4.2 GPU Cost Comparison
In this section, we report the maximum GPU memory allocated during the inference across both datasets for each method mentioned in Section 4.1. The results are shown in Table 1. From the results, we can find that M+ has the lowest GPU memory cost except for Llama-3.1-8B-16k. The reason that M+ uses fewer tokens but costs more GPU is that we have 12,800 tokens in each layer, while Llama-3.1-8B-16k has only one layer of 16k tokens. Therefore, we propose to offload the memory tokens on CPU, and reload them into GPU when the calculation reaches a certain layer. By “CPU offloading”, we specifically mean offloading the memory vectors present in each layer of the model. It is important to note that other models, such as Llama-3.1-8B do not have memory vectors, so our CPU offloading can only be applied to MemoryLLM and M+. With that, we can sacrifice some I/O time cost but substantially decrease the GPU cost without affecting the performance. This leads to “M+ (offload)” which achieves the least GPU memory consumption. We also include the GPU cost of MemoryLLM-8B, which is the model obtained after Stage 1 described in Section 3.2.4. This shows that the Long-Term Memory does not incur more GPU costs.
Table 1: GPU Memory Cost Comparison. MemoryLLM-8B is the model obtained after Stage 1 training, serving as an ablation study.
@@ 4.4 Experimental Results on (Relatively) Short Documents
We evaluate the performance of M+ and Llama-3.1-8B on relatively short documents using the LongBench benchmark, considering input lengths of 8k and 16k tokens. The evaluation metric is QA-F1, following Bai et al. (2023). The results, presented in Table 2, show that M+ could match the performance of Llama-3.1-8B on 4 out of 6 datasets, except for hotpotqa and musique. This performance difference can be attributed to two primary factors: (1) Random Dropping Mechanism: M+ employs a random dropping mechanism that can lead to partial information loss. For instance, processing an 8k input requires splitting it into 12 chunks (each chunk being 512 tokens), while the last 2k tokens are directly included in the generation context window. The first 6k tokens (12 chunks) are compressed into $2755.6256+256*\frac{N-K}{N}+\cdots+256*(\frac{N-K}{N})^{11}=$ 2755.6 tokens (with around 316.4 memory tokens dropped). For 16k tokens, the first 14k tokens are compressed 5530 tokens (with around 1638 tokens dropped). As some tokens are dropped, the performance may get affected slightly. Note that this could lead to a longer context window while sacrificing some of the performance in relatively shorter context tasks. (2) Limited Cross-Chunk Attention: When processing chunks into memory, M+ uses the last $K$ tokens and the hidden states from the new chunk as input (illustrated in Figur e1). In contrast, Llama-3.1-8B processes each chunk with access to all previous tokens, enabling cross-attention between chunks. While this approach allows Llama-3.1-8B to maintain full attention across chunks, it comes with significantly higher computational and memory costs due to the quadratic scaling of transformer computations. In comparison, M+ achieves linear computational scaling, making it more GPU-memory-efficient for processing extremely long inputs (see Section 4.2), albeit with some trade-off in performance in tasks with relatively shorter contexts.
@@ 4.5.1 Ablation Study on long-term memory
Long-term memory does not affect the performance on relatively short documents
To show whether long-term memory affects the model’s performances on relatively short documents, we conduct ablation study with models from three different stages on the dataset LongBench while fixing the context window as 8k. The results are shown in Table 3. From the table we can see that M+ has similar performance as MemoryLLM-8B-Long on 8k context window. This means adding long-term memory does not affect the performance on relatively short documents. Meanwhile, MemoryLLM-8B-Long is significantly better than MemoryLLM-8B, which can be attributed to the inclusion of longer training examples in Stage 2 (4k–64k), whereas fineweb-edu (used in Stage 1) contains very few examples longer than 4k.
Table 3: Ablation study of the effects of different stages on LongBench.
2wikimqa hotpotqa qasper musique multifieldqa_en narrativeqa Avg
MemoryLLM-8B (Stage 1, 8k) 32.30 33.39 23.88 12.37 35.91 21.46 26.55
MemoryLLM-8B-Long (Stage 2, 8k) 32.23 37.86 31.62 20.35 42.16 23.49 31.29
M+ (Stage 3, 8k) 33.12 37.99 29.91 20.68 40.11 24.18 31.00
@@ 4.6 Analysis
4.6.1 Model Quality within Context Window
M+ uses 12,800 memory tokens alongside a 2,048-token generation context window. In this section, we evaluate the model’s performance within the standard 2,048-token context window to ensure that the addition of memory does not degrade its base capability. We randomly select 1,000 examples from the fineweb-edu dataset (snapshot CC-MAIN-2024-10), which does not overlap with the training data. For this evaluation, we cap the input sequence length at 2,048 tokens and report perplexity for both M+ and LLaMA-3.1-8B. The results show that LLaMA-3.1-8B achieves a perplexity of 1.9734, while M+ records a similar perplexity of 1.9828. These findings indicate that M+ maintains competitive performance on documents shorter than 2,048 tokens, confirming that the base model’s quality within the context window remains intact.
@@ 4.6.2 Retrieval Quality
In our implementation, the long-term memory is initially of size 5120, and then it gradually increases to 80k in our knowledge retention experiments (it hits 81,276 when there 160k tokens are injected). To access retrieval quality, we leverage the knowledge retention task with SQuAD dataset, where the first $K=256$ tokens are critical for answering the questions. These $K=256$ tokens are denoted as ground-truth tokens. We track the number of ground-truth tokens in the long-term memory and how many tokens are retrieved back into the “Extracted LTM” pool in Figure 1 when queried after various numbers of tokens are injected. We present the results in Figure 6, demonstrating the retrieval quality as more tokens are dropped from the memory pool to the long-term memory. From the figure we can see that around 30% tokens are retrieved. For reference, random retrieval would retrieve $2,560/81,276=3\%$ tokens.